OpenStudy (anonymous):

Two numbers whose product is 1 are called ______. equals fractions opposites reciprocals

pooja195 (pooja195):

Well this one is kind of hard to explain but think about the key word SAME...

OpenStudy (anonymous):

it is not a fraction nor recipracle because they have to do with fractions

OpenStudy (anonymous):

1 and 1 is the same so it would not be opp so the answer is equals @minecraftdude1524 @pooja195

OpenStudy (anonymous):

ii explaind it leaad him through the dark forest @pooja195

pooja195 (pooja195):

Yeah well you still gave away the answer which is not cool.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

and no , it is not "equals"

OpenStudy (anonymous):

?????

OpenStudy (texaschic101):

reciprocals... examples : 2/3 * 3/2 = 6/6 = 1 1/2 * 2/1 = 2/2 = 1 4/3 * 3/4 = 12/12 = 1